Circular No. 7661

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 N2 (LINEAR) An apparently asteroidal object reported by LINEAR (discovery ob-

servation below), which was posted on the NEO Confirmation Page, has been found to have a diffuse coma and a faint 5′′ tail in p.a. ∼ 90^{o} in a 3 min r-band CCD exposure taken with the 0.6-m reflector at Table Mountain by M. Hicks, D. Esqueda, and T. Ha.

2001 UT  α_{2000}  δ_{2000}  m_{2}
July 11.38017  23^{h}03^{m}09.^{s}94  +39^{o}05′57′′.8  18.1

Additional astrometry and the following preliminary parabolic orbital elements from 18 observations, July 11–13, appear on MPEC 2001-N29:

T = 2002 Aug. 29.527 TT  ω = 138.^{o}456
Ω = 48.346 }2000.0
q = 3.17991 AU  i = 136.969

SUPERNOVA 2001db IN NGC 3256 R. Maiolino and M. Della Valle, Osservatorio Astrofisico di Arcetri,

Florence; L. Vanzi, European Southern Observatory (ESO); and F. Mannucci, Consiglio Nazionale delle Ricerce, Florence, report the discovery of a supernova during a campaign of near-infrared monitoring of luminous infrared galaxies: “SN 2001db was discovered at mag 16.27 in a K_{s}-band image obtained on 2001 Jan. 9.3 UT with the ESO New Technology Telescope (+ SOFI), when compared with a previous archival K_{s} image from Nov. 1999. To our knowledge, this is the first supernova ever discovered in the infrared. SN 2001db is located at α = 10^{h}27^{m}50^{s}.4, δ = −43^{o}54′21′′ (equinox 2000.0), which is 6′′ west and 6′′ south of the K_{s} nucleus of NGC 3256 (uncertainty on relative position ∼ 0′′.4). Two additional K_{s}-band images taken on Feb. 9.25 and Mar. 8.1 show that SN 2001db faded to K_{s} = 17.11 and 17.55, respectively. Infrared and optical spectra obtained at the ESO Very Large Telescope on Apr. 21.1 (with the Infrared Spectrometer And Array Camera) and May 17.0 (with the Focal Reducer/low-dispersion Spectrograph) show a broad component (∼ 5000 km/s FWHM) of the Paβ and of Hα flanked by faint P-Cyg absorptions, thereby suggesting this to be a type-II event caught ∼ 5 months after maximum. This implies that our first detection of the supernova (on Jan. 9) was ∼ 1 month after maximum light.”

2001 July 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green